TERAHERTZ SPECTROSCOPY AND GLOBAL ANALYSIS OF H_{3}O^{+}

\underline{SHANSHAN YU}, BRIAN J. DROUIN, JOHN C. PEARSON AND HERBERT M. PICKETT, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

Eight ground state inversion transitions of H_{3}O^{+} in the 0.9-1.6 THz region have been measured using the frequency multiplier submillimeter spectrometer (FMSS) at JPL. The accuracy of these measurements is estimated to be 300 kHz. The ions were generated in a DC discharge through a gas mixture of a few mTorr of H_{2} and 30 mTorr of H_{2}O. A multistate analysis was carried out for H_{3}O^{+}, which includes the lines observed in this work, previous submillimeter and terahertz inversion transitions in the ground state, and previous infrared data on all the four vibrational fundamental bands. Accurate molecular parameters were obtained with taking into account the strong Coriolis interaction between the symmetric OH stretching mode ν_{1} and the doubly degenerate asymmetric OH stretching mode ν_{3}. Frequency predictions, particularly for high-J transitions in the ground state and in the ν_{1} and ν_{3} fundamental bands, have been greatly improved by including in our analysis 200 more high-J transitions, which were excluded in previous analyses. The more precise measurements and new predictions reported here will support the analysis of astronomical observations by the future high resolution spectroscopy telescopes such as Herschel, SOFIA, and ALMA.